OpenStudy (timaashorty):

Please help ! Solve the quadratic equation by completing the square -3x^2+9x=1

OpenStudy (anonymous):

\[-\left( 3x2-9x-\frac{ 27}{ 4 } \right)\]

OpenStudy (timaashorty):

The answer choices : A. \[\frac{2 }{3 } \pm \sqrt\frac{ 6 }{6 } \] B.\[-3\pm \sqrt\frac{ 69}{3 } \] C.\[3\pm \sqrt\frac{ -3 }{3 } \] D.\[\frac{ 3 }{2 } \pm \sqrt\frac{ 69 }{6 } \]

OpenStudy (anonymous):

Oh, D

OpenStudy (timaashorty):

Thank you!

OpenStudy (anonymous):

No prob

OpenStudy (timaashorty):

But are you 100% sure ? lol

OpenStudy (anonymous):

110%

OpenStudy (timaashorty):

Haha Thanks Again Then (: